OpenStudy (anonymous):

Which expressions have a value of –23? Choose all answers that are correct. A. 11 – 34 B. –9 – (–14) C. 25 – 2 D. –11 – 12

OpenStudy (anonymous):

B and D?

OpenStudy (ryanwhite499):

11 - 34 is what on a calculator

OpenStudy (anonymous):

23

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Lol i read it wrong

OpenStudy (ryanwhite499):

Yes... 11 - 34 is...

OpenStudy (anonymous):

11 minus 34 not 34 - 11

OpenStudy (anonymous):

its -23

OpenStudy (ryanwhite499):

Correct

OpenStudy (anonymous):

A B and D?

OpenStudy (ryanwhite499):

So a is one

OpenStudy (ryanwhite499):

Lets check

OpenStudy (ryanwhite499):

B is not on them see −9−(−14) =−9−(−14) =−9+14 =5

OpenStudy (anonymous):

-9 - (-14) 9 + 14 23 -23

OpenStudy (anonymous):

i mean -9 + 14

OpenStudy (ryanwhite499):

C does not define the answer either because it is not a negative

OpenStudy (texaschic101):

no wait......it is just A and D

OpenStudy (ryanwhite499):

Yes that's right... A and D is correct

OpenStudy (anonymous):

last question

OpenStudy (anonymous):

i always forget

OpenStudy (anonymous):

-5 + -5 = -10 -5 - (-5) = -10 right?

OpenStudy (ryanwhite499):

Its True

OpenStudy (anonymous):

ok i always thought that they where different

OpenStudy (ryanwhite499):

Your equation is perfect

OpenStudy (ryanwhite499):

−5−5 =−5+−5 =−10

OpenStudy (texaschic101):

no... -5 + (-5)......neg times a pos is a neg -5 - 5 = -10

OpenStudy (ryanwhite499):

Yes thats what she said

OpenStudy (ryanwhite499):

Its -10

OpenStudy (texaschic101):

yes but she put -5 - (-5)....that would change it

OpenStudy (ryanwhite499):

Yes it will

OpenStudy (texaschic101):

but the answer is -10

OpenStudy (ryanwhite499):

Make sure you put the correct equation so we can help you answer correctly